OpenStudy (anonymous):

What volume of 0.155M Na3PO4 solution is necessary to completely react with 90.7mL of 0.102M CuCl2?

OpenStudy (aaronq):

start by writing (and balancing) the equation for the reaction

OpenStudy (anonymous):

\[2Na _{3}PO _{4}+3CuCl _{2}\rightarrow6NaCl+Cu _{3}(PO _{4})_{2}\]

OpenStudy (anonymous):

So I would start off with .00925 mol of\[CuCl _{2}\]

OpenStudy (aaronq):

for problems involving stoichiometry, you have to convert whatever information they give you to moles (signified by the symbol "n") convert 90.7mL of 0.102M \(CuCl_2\) to moles Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients, \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) once you find the moles of \(Na_2PO_4\), you can convert back to volume using the concentration given.

OpenStudy (anonymous):

Does 25.13L sound right?

OpenStudy (aaronq):

no thats way too much

OpenStudy (anonymous):

I got .0061676 mol of NaPO4

OpenStudy (anonymous):

then divided the molarity given by the previous number (moles)

OpenStudy (aaronq):

and you got 25 Litres?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (aaronq):

\(\dfrac{0.0907\;L*0.102M}{3}=\dfrac{n_{Na_3PO4}}{2}\) \(n_{Na_3PO_4}=0.0604666666666667\) \(L_{solution}=\dfrac{n_{Na_3PO_4}}{M_{Na_3PO_4}}=\dfrac{0.0604666666666667}{0.155}\) =0.39010 L

OpenStudy (anonymous):

Oh... I completely flipped the moles and Molarity. *facepalm* Makes more sense now! Thanks!

OpenStudy (aaronq):

no problem, hope its clear !